OpenStudy (anonymous):

Write the equation of the line that passes through (2, 5) with a slope of 3 in point-slope form. Answer y – 5 = 3(x – 2) y + 8 = 3(x + 5) y – 2 = 3(x – 5) y + 5 = 3(x + 2)

jimthompson5910 (jim_thompson5910):

y-y1 = m(x-x1) y-5 = 3(x-2) So the answer is choice A